# An ALMA spectroscopic survey of the Planck high-redshift object PLCK G073.457.5 confirms two protoclusters

### v1

@@ 5.2 Depletion timescales
We also show for comparison the values of the CO-detected cluster and protocluster galaxies described in Sect. 4.1. The gas masses from the literature have been derived from various CO transitions and values – e.g., in Noble et al. (2017, 2019); Rudnick et al. (2017), in Ivison et al. (2013) – and in some cases a stellar mass-dependent correction was applied resulting in final values ranging from 4.16 to 6.09 (Hayashi et al., 2018).
The conversion factor adopted by Sargent et al. (2014) varies with the metallicity following Wolfire et al. (2010) and is 4.4 at solar metallicity and always above 3. The gas masses in Liu et al. (2019) are derived from the continuum. We corrected all the gas mass estimates from this literature cluster sample to have the same conversion factor of 4.36 (when possible), which is the value we used for G073. The scaling relations used to represent normal star-forming galaxies in the field are consistent with the adopted conversion.
We find that the depletion timescales of the G073 CO emitters are scattered around the values observed in normal SFGs (). The median depletion times for the G073 sources and for the cluster members are () Gyr and () Gyr, respectively, thus we find no significant difference.
@@ 6.1 The large-scale structure probed by G073
For the simulation comparison, Gouin et al. (2022) chose the redshift snapshot closest to the actual redshifts of the G073 structures, selected 30 of the most star-forming halos (defined to include all gas cells in a friends-of-friends group), then accounted for observational biases by including galaxies within a cylinder (centered on a given star-forming halo) with a diameter of 2.4′and a length encompassing in redshift for the structure, and in redshift for the structure. Lastly, an SFR cut of 30 M⊙ yr-1 was applied to the galaxies within the simulated cylinders. This is comparable to our observational sensitivity – our 5 line luminosity threshold is about 0.3 Jy km s-1, corresponding to a gas mass of M⊙ (assuming CO(3–2) at ) or a SFR of 30 M⊙ yr-1 using the scaling relation from Sargent et al. (2014). For Ci(1–0) at , this limit corresponds to a gas mass of M⊙, or a SFR of 100 M⊙ yr-1.
@@
In Fig. 8 we reproduce the panels in figures 6 and 7 from Gouin et al. (2022) relating to the and structures in G073; we show the total SFR versus total stellar mass (i.e., the sum of all the galaxies within the cylinder) of all 30 simulated structures, and the total number of galaxies versus total stellar mass for the same simulated sample. We also show the same quantities for the observed galaxies as reported in Kneissl et al. (2019). Gouin et al. (2022) selected one of the 30 simulated structures that best matched the observations (highlighted in red), and we show individual SFRs and stellar masses of these specific simulated structures compared to the galaxies from Kneissl et al. (2019). Gouin et al. (2022) conclude that the TNG300 simulation is able to reproduce the PHz selection, and that the simulated structures are comparable to the observations.
@@
With our new ALMA observations spectroscopically confirming six galaxies around and five galaxies around , we would like to ask if the conclusions from Gouin et al. (2022) concerning G073 have changed. In Fig. 8 we therefore update the observed data points with values derived in this work. We find that the total observed SFR within the structure increases while the total stellar mass decreases, although not substantially, while the structure is in good agreement. The properties of the individual galaxies within G073 also remain consistent with the simulated structures selected by Gouin et al. (2022) to be the most similar to the observations from Kneissl et al. (2019).
@@
Regarding the fates of the most star-forming halos found in the TNG300 simulation, Gouin et al. (2022) found that 60–70% of the simulated structures (which spanned a redshift range of 1.3 to 3.0) evolved into M⊙ galaxy clusters by redshift 0. Moreover, they found that protoclusters with more star-forming galaxies and with that star formation evenly spread across the galaxies (i.e., systems without a single extremely luminous galaxy alongside a number of faint galaxies) are more likely to become clusters by . This is roughly in line with our observations, so we conclude that the results from Gouin et al. (2022) suggest a high probability that both structures in G073 will ultimately evolve into a galaxy clusters.
@@
While these values are nowhere near the mass of a redshift 0 galaxy cluster, it is worth emphasizing that our line luminosity limits are modest. For example, van der Burg et al. (2013) fit Schechter functions to a sample of 10 spectroscopically-confirmed galaxy clusters around , splitting the sources into a sample of star-forming galaxies and a sample of quiescent galaxies. Using their fit to the star-forming galaxies, and assuming we are only sensitive to galaxies with stellar masses M⊙, we find that we would expect to find about 4 times more galaxies with stellar masses M⊙ than with stellar masses M⊙ (corresponding to a total of about 25 galaxies in each system). This is again not enough mass to account for 10M⊙, but growth through accretion outside of the field-of-view is expected.
@@
To investigate possible environmental effects on the gas properties of cluster and protocluster galaxies, we derive the field-relative gas depletion timescales and gas-to-stellar mass ratios by dividing the measured values from our sample and the collection from the literature by those of coeval field galaxies obtained through the Liu et al. (2019) scaling relations. The predicted field values are calculated using the redshift and the stellar mass of each source and ignoring corrections related to offsets to the main sequence. The broad range of stellar masses (about (0.5–17) M⊙ or 1.5 dex span) and of SFRs (about 3–600 M⊙ yr-1, or 2.3 dex span) of the G073 sources, combined with the cluster members, allows us to investigate possible dependencies with these parameters. The properties relative to field galaxies are shown as a function of SFR and stellar mass in Fig. 9.
We find that 51% of the cluster and protocluster members and 30% of the G073 sources have gas-to-stellar mass ratios that are consistent with coeval field galaxies, while 39% and 70% have larger ratios, respectively. This is consistent with previous results which find that CO emitters in high- clusters have field-like or enhanced molecular gas fractions (i.e., Noble et al., 2017; Rudnick et al., 2017; Hayashi et al., 2018). Interestingly, the excess in gas fraction is observed at all masses, but there seems to be a mild trend with SFR. Although these results might suffer from a bias in favour of gas rich galaxies, it is expected that high-density environments at high redshift favour cold gas buildup, and the systems with the highest gas fractions are also the most star-forming ones.
@@ 6.3 The nature of 850 m-bright sources in PHz fields
It is also worth discussing the fact that no SCUBA-2-selected sources appear to be in the redshift 1.5 or 2 structures. These sources all have mJy (and three have mJy), and were part of a larger study of 61 PHz fields followed up by SCUBA-2 (MacKenzie et al., 2017). Of these 61 fields, 51 were protocluster candidates (so not strong gravitational lenses), and ultimately contained a factor of 6 enhancement in number density for sources above 8 mJy compared to blank-field surveys. This implied that the PHz objects at least contain projected overdensities of sources at 850 m on the sky. MacKenzie et al. (2017) also fitted for photometric redshifts using the three Herschel-SPIRE bands and photometry at 850 m from SCUBA-2, specifically finding photometric redshifts for the four sources followed up in our programme (although with large uncertainties) – effectively consistent with one of these sources being at .
@@ 7 Summary and conclusions
We have used ALMA to search for and protocluster members in the Planck-selected field G073. In eight pointings targeting bright and red Herschel-SPIRE sources and four pointings targeting high S/N SCUBA-2 sources across the field, we have identified 27 individual galaxies through a combination of continuum and line emission, confirming six spectroscopic redshifts around and five spectroscopic redshifts around . We have also found that the SCUBA-2 sources are not part of either of these protoclusters, instead likely lying in a more distant structure (or structures) at .
@@
By fitting near-IR (m) and far-IR (up to 2 mm) photometry to our sources, we derived star-formation rates and stellar masses, and we have also estimated gas masses using the CO and Ci line detections from our ALMA programme. Four galaxies are considered starbursts (having SFRs a factor of 4 larger than the MS), and the average gas-to-stellar mass ratio () of our protocluster galaxies is higher than typical scaling relations for field galaxies.
@@
We have compared our more detailed view of G073 to the TNG300 simulation by updating the results from Gouin et al. (2022) with spectroscopic redshift information. Our observations remain consistent with the simulated star-forming halos at high redshift selected from the simulation. Since about 60–70% of the simulated halos collapse into galaxy clusters by redshift zero, we conclude that the two structures uncovered in G073 are comparably likely to become clusters as well. We estimate the total dark matter halo mass in observed galaxies to be around in each structure, although we speculate that there are still a factor of 4 more galaxies with stellar masses an order of magnitude below our detection threshold of M⊙.
@@
We show that the two structures spectroscopically detected by our observations are separated by about 3′in projection on the sky. Relative to Planck’s 5′beam originally used to select this field, this is in-line with previous claims that at least some of the PHz objects are line-of-sight alignments (e.g. Flores-Cacho et al., 2016), and indeed the presence of an overdensity of SCUBA-2 sources at higher redshift could indicate that this field contains three or more protoclusters. While this poses an observational challenge for future follow-up studies of a statistically significant sample of PHz objects, it also provides the potential for extracting more information than achievable from typical single-system fields. For example, the upcoming ALMA receiver upgrade (Carpenter et al., 2023) will boost the facility’s spectral coverage by a factor of 4, meaning that with short observations of single PHz fields (such as G073) we can spectroscopically identify several protoclusters. A modest observing programme targeting of order 10 PHz objects therefore has the potential to find dozens of protoclusters, easily providing powerful population statistics of a homogeneously-selected sample of star-forming protoclusters (and their embedded galaxies) around the peak of the star formation activity of the Universe.
@@ Acknowledgements.
All of the data presented in this paper are publicly available. The ALMA data can be found under the two programmes 2013.1.01173.S (PI: R. Kneissl) and 2018.1.00562.S (PI: R. Hill). The Spitzer data are from Program ID 11004 (PI: H. Dole) and the CFHT data are from Proposal IDs 13BF12 and 14BF08 (PI: H. Dole).
@@ Appendix A ALMA spectra
Appendix B Near-IR photometric data
Here we provide the near-IR photometry used to fit SEDs and derive physical properties of the galaxies in G073. The and -band data are from WIRcam on the CFHT, and the 3.6 m and 4.5 m data are from Spitzer-IRAC.

### v2

@@ 5.2 Depletion timescales
We also show for comparison the values of the CO-detected cluster and protocluster galaxies described in Sect. 4.1. The gas masses from the literature have been derived from various CO transitions and values – for example in Noble et al. (2017), Noble et al. (2019), and Rudnick et al. (2017) and in Ivison et al. (2013) – and in some cases a stellar mass-dependent correction was applied resulting in final values ranging from 4.16 to 6.09 (Hayashi et al. 2018).
The conversion factor adopted by Sargent et al. (2014) varies with the metallicity following Wolfire et al. (2010) and is 4.4 at solar metallicity and always above 3. The gas masses in Liu et al. (2019) are derived from the continuum. We corrected all the gas mass estimates from this literature cluster sample to have the same conversion factor of 4.36 (when possible), which is the value we used for G073. The scaling relations used to represent normal SFGs in the field are consistent with the adopted conversion.
We find that the depletion timescales of the G073 CO emitters are scattered around the values observed in normal SFGs (). The median depletion times for the G073 sources and for the cluster members are () Gyr and () Gyr, respectively; thus, we find no significant difference.
@@ 6.1 The large-scale structure probed by G073
For the simulation comparison, Gouin et al. (2022) chose the redshift snapshot closest to the actual redshifts of the G073 structures, selected 30 of the most star-forming halos (defined to include all gas cells in a friends-of-friends group), then accounted for observational biases by including galaxies within a cylinder (centred on a given star-forming halo) with a diameter of 2.4′and a length encompassing in redshift for the structure, and in redshift for the structure. Lastly, an SFR cut of 30 M⊙ yr-1 was applied to the galaxies within the simulated cylinders. This is comparable to our observational sensitivity – our 5 line luminosity threshold is about 0.3 Jy km s-1, corresponding to a gas mass of M⊙ (assuming CO(3–2) at ) or a SFR of 30 M⊙ yr-1 using the scaling relation from Sargent et al. (2014). For Ci(1–0) at , this limit corresponds to a gas mass of M⊙, or a SFR of 100 M⊙ yr-1.
@@
In Fig. 8 we reproduce the panels in figures 6 and 7 from Gouin et al. (2022) relating to the and structures in G073; we show the total SFR versus total stellar mass (i.e., the sum of all the galaxies within the cylinder) of all 30 simulated structures, and the total number of galaxies versus total stellar mass for the same simulated sample. We also show the same quantities for the observed galaxies as reported in Kneissl et al. (2019). Gouin et al. (2022) selected one of the 30 simulated structures that most closely matched the observed total stellar mass, SFR, and number of SFGs (highlighted in red), and we show individual SFRs and stellar masses of these specific simulated structures compared to the galaxies from Kneissl et al. (2019). Gouin et al. (2022) conclude that the TNG300 simulation is able to reproduce the PHz selection, and that the simulated structures are comparable to the observations.
@@
With our new ALMA observations spectroscopically confirming six galaxies around and five galaxies around , we would like to ask if the conclusions from Gouin et al. (2022) concerning G073 have changed. In Fig. 8 we therefore update the observed data points with values derived in this work. We find that the total observed SFR within the structure increases while the total stellar mass decreases, although not substantially, while the structure is in good agreement. The properties of the individual galaxies within G073 also remain consistent with the simulated structures selected by Gouin et al. (2022) to be the most similar to the observations from Kneissl et al. (2019). It is worth noting that while there still appear to be some discrepancies between the simulated protocluster galaxies and the observations (e.g. higher observed stellar masses and SFRs compared to the selected simulation at ), deciding which simulation to focus on for the comparison is subjective; a more statistically robust comparison would marginalize over all of the simulated protoclusters, and so we emphasize here that this comparison only qualitatively illustrates broad agreement.
@@
Regarding the fates of the most star-forming halos found in the TNG300 simulation, Gouin et al. (2022) found that 60–70% of the simulated structures (which spanned a redshift range of 1.3 to 3.0) evolved into M⊙ galaxy clusters by redshift 0. Moreover, they found that protoclusters with more SFGs and with that star formation evenly spread across the galaxies (i.e., systems without a single extremely luminous galaxy alongside a number of faint galaxies) are more likely to become clusters by . This is roughly in line with our observations, so we conclude that the results from Gouin et al. (2022) suggest a high probability that both structures in G073 will ultimately evolve into a galaxy clusters.
@@
While these values are nowhere near the mass of a redshift 0 galaxy cluster, it is worth emphasizing that our line luminosity limits are modest. For example, van der Burg et al. (2013) fit Schechter functions to a sample of 10 spectroscopically confirmed galaxy clusters around , splitting the sources into a sample of SFGs and a sample of quiescent galaxies. Using their fit to the SFGs, and assuming we are only sensitive to galaxies with stellar masses M⊙, we find that we would expect to find about 4 times more galaxies with stellar masses M⊙ than with stellar masses M⊙ (corresponding to a total of about 25 galaxies in each system). This is again not enough mass to account for 10M⊙, but growth through accretion outside of the field-of-view is expected.
@@
To investigate possible environmental effects on the gas properties of cluster and protocluster galaxies, we derived the field-relative gas depletion timescales and gas-to-stellar mass ratios by dividing the measured values from our sample and the collection from the literature by those of coeval field galaxies obtained through the Liu et al. (2019) scaling relations. The predicted field values are calculated using the redshift and the stellar mass of each source and ignoring corrections related to offsets to the MS. The broad range of stellar masses (about (0.5–17) M⊙ or 1.5 dex span) and of SFRs (about 3–600 M⊙ yr-1, or 2.3 dex span) of the G073 sources, combined with the cluster members, allows us to investigate possible dependences with these parameters. The properties relative to field galaxies are shown as a function of SFR and stellar mass in Fig. 9.
We find that 51% of the cluster and protocluster members and 30% of the G073 sources have gas-to-stellar mass ratios that are consistent with coeval field galaxies, while 39% and 70% have higher ratios, respectively. This is consistent with previous results that find that CO emitters in high- clusters have field-like or enhanced molecular gas fractions (i.e., Noble et al. 2017; Rudnick et al. 2017; Hayashi et al. 2018). Interestingly, the excess in gas fraction is observed at all masses, but there seems to be a mild trend with SFR. Although these results might suffer from a bias in favour of gas rich galaxies, it is expected that high-density environments at high redshift favour cold gas buildup, and the systems with the highest gas fractions are also the most star-forming ones.
@@ 6.3 The nature of 850 m-bright sources in PHz fields
It is also worth discussing the fact that no SCUBA-2-selected sources appear to be in the redshift 1.5 or 2 structures. These sources all have mJy (and three have mJy), and were part of a larger study of 61 PHz fields followed up by SCUBA-2 (MacKenzie et al. 2017). Of these 61 fields, 51 were protocluster candidates (so not strong gravitational lenses), and ultimately contained a factor of 6 enhancement in number density for sources above 8 mJy compared to blank-field surveys. This implied that the PHz objects at least contain projected overdensities of sources at 850 m on the sky. MacKenzie et al. (2017) also fitted for photometric redshifts using the three Herschel-SPIRE bands and photometry at 850 m from SCUBA-2, specifically finding photometric redshifts for the four sources followed up in our programme (although with large uncertainties) – effectively consistent with one of these sources being at .
@@
6.4 The nature of the Planck high- sources
The PHz sources are the brightest 5′-scale submm sources on the sky. By selection, their thermal submm SEDs peak between 350 and 850 m, corresponding to redshifts of around 1.5–4, and their brightnesses (if unlensed) imply SFRs 25,000 M⊙ yr-1 (Planck Collaboration XXXIX 2016). We now know that only a few percent are strong lenses (e.g. Cañameras et al. 2015), with the remaining objects consisting of groups of submm sources that exhibit extended emission that is absent in randomly selected Herschel samples (e.g. Lammers et al. 2022).
The PHz selection clearly favours fields where submm sources are aligned along the line of sight, but the physical associations (i.e. the redshift distributions) remain debated. Nonetheless, observations are routinely identifying two or more galaxy groups along the line of sight (PLCK G95.561.6 at and (Flores-Cacho et al. 2016); PLCK G237.0142.50 at and (Polletta et al. 2021); and PLCK G073.457.5 at and , as confirmed here and initially found by Kneissl et al. 2019), or a single galaxy group with numerous galaxies in the foreground and background (PLCK G176.60+59.01 at ; PLCK DU G059.137.4 at ; PLCK DU G124.168.8 at ; and PHz G191.2462.04 at – see Polletta et al. 2022, 2024).
Ultimately, based on the limited and biased spectroscopic follow-up programmes that have been carried out so far, there seems to be evidence that the Planck objects consist of sometimes one, but often more, structures with redshifts between about 2 and 3, each traced by several Herschel sources. There could of course be an observational bias against fields containing no protoclusters (meaning that all of the line of sight galaxies are at different redshifts) as securing the redshifts of numerous unassociated galaxies is challenging, but simulations seem to suggest otherwise (e.g., Negrello et al. 2017; Gouin et al. 2022). However the question still remains as to how massive these structures are, requiring deep spectroscopic observations sensitive to both dusty obscured sources and optically luminous sources over wide areas.
@@ 7 Summary and conclusions
We used ALMA to search for and protocluster members in the Planck-selected field G073. In eight pointings targeting bright and red Herschel-SPIRE sources and four pointings targeting high S/N SCUBA-2 sources across the field, we have identified 27 individual galaxies through a combination of continuum and line emission, confirming six spectroscopic redshifts around and five spectroscopic redshifts around . We have also found that the SCUBA-2 sources are not part of either of these protoclusters; they likely lie in a more distant structure (or structures) at .
@@
By fitting NIR (m) and FIR (up to 2 mm) photometry to our sources, we derived SFRs and stellar masses, and we estimated gas masses using the CO and Ci line detections from our ALMA programme. Four galaxies are considered starbursts (having SFRs a factor of 4 higher than the MS), and the average gas-to-stellar mass ratio () of our protocluster galaxies is higher than typical scaling relations for field galaxies.
@@
We compared our more detailed view of G073 to the TNG300 simulation by updating the results from Gouin et al. (2022) with spectroscopic redshift information. Our observations remain consistent with the simulated star-forming halos at high redshifts selected from the simulation. Since about 60–70% of the simulated halos collapse into galaxy clusters by redshift zero, we conclude that the two structures uncovered in G073 are comparably likely to become clusters as well. We estimate the total dark matter halo mass in observed galaxies to be around in each structure, although we speculate that there are still a factor of 4 more galaxies with stellar masses an order of magnitude below our detection threshold of M⊙.
@@
We show that the two structures spectroscopically detected by our observations are separated by about 3′in projection on the sky. Relative to Planck’s 5′beam originally used to select this field, this is in line with previous claims that at least some of the PHz objects are line-of-sight alignments (e.g. Flores-Cacho et al. 2016), and indeed the presence of an overdensity of SCUBA-2 sources at higher redshifts could indicate that this field contains three or more protoclusters. While this poses an observational challenge for future follow-up studies of a statistically significant sample of PHz objects, it also provides the potential for extracting more information than achievable from typical single-system fields. For example, the upcoming ALMA receiver upgrade (Carpenter et al. 2023) will boost the facility’s spectral coverage by a factor of 4, meaning that with short observations of single PHz fields (such as G073) we will be able to spectroscopically identify several protoclusters. A modest observing programme targeting of order ten PHz objects therefore has the potential to find dozens of protoclusters, easily providing powerful population statistics of a homogeneously selected sample of star-forming protoclusters (and their embedded galaxies) around the peak of the star formation activity of the Universe.
@@ Acknowledgements.
All of the data presented in this paper are publicly available. The ALMA data can be found under the two programmes 2013.1.01173.S (PI: R. Kneissl) and 2018.1.00562.S (PI: R. Hill). The Spitzer data are from programme ID 11004 (PI: H. Dole) and the CFHT data are from Proposal IDs 13BF12 and 14BF08 (PI: H. Dole).
@@ Appendix A ALMA spectra
Appendix B NIR photometric data
Here we provide the NIR photometry used to fit SEDs and derive physical properties of the galaxies in G073. The and -band data are from WIRcam on the CFHT, and the 3.6 m and 4.5 m data are from Spitzer-IRAC.